OpenStudy (anonymous):

Given the set of inputs {-4,-1,0,1,4} determine if the following relation is a function. if not explain. 1) 2x+7 ...what am I suppouse to do?

OpenStudy (anonymous):

You can plug each of those values into the equation and make sure that no one x value has more than one y-value. The simple answer is that it can't since the function is linear.

OpenStudy (zzr0ck3r):

you must show that if x = y then f(x) = f(y) so ... x = y 2x=2y 2x+7=2y+7 f(x)=f(y)

OpenStudy (zzr0ck3r):

and you must show that all of the domain(inputs) map to somewhere . In other words it is ok to plug in all the inputs. And in this case it is.

OpenStudy (anonymous):

So is 2x+7 a function or not?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

What about |x|? How will that work with the inputs?

OpenStudy (anonymous):

For any conceivable value of x that you put in, do you get out exactly one y?

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

Well, can you think of ANY value of x that will give you 2 numbers back? You can also look at it graphically. This is what |x| looks like: 

OpenStudy (anonymous):

When you start seeing this, you're not in function territory: